OpenStudy (anonymous):

A druggist wishes to select three brands of aspirin to sell in his store. He has five major brands to choose from: A, B, C, D and E. If he selects the three brands at random, what is the probability that at least one of the two brands B and C

OpenStudy (anonymous):

Solve these problems by finding out probability of failure and subtracting it from 1 P(fail) = (3/5)(3/5)(3/5) chose ADE out of the five each time = 0.216 P(success) = 1 - P(fail) = 0.784

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i think that might not be right

OpenStudy (anonymous):

ya i just tried and its wrong :(

OpenStudy (anonymous):

there are a couple ways to do it the idea of computing 1 minus the probability than none are B and C is a good one

OpenStudy (anonymous):

that would be \[\frac{3}{5}\times \frac{2}{4}\times \frac{1}{3}\] however

OpenStudy (anonymous):

I am assuming there are plenty of each and that he makes three withdrawals, missing B and C each time, thus failing three times in a row. Success is 1 - failure probability, I believe.

OpenStudy (anonymous):

then your answer is \[1-\frac{3}{5}\times \frac{2}{4}\times \frac{1}{3}\] whatever that is

OpenStudy (anonymous):

Five brands (lots of each) seemed different from five objects, A to E.

OpenStudy (anonymous):

thanks!!! that was right

OpenStudy (anonymous):

He has five major brands to choose from not he is going to pick bottle at random from an unlimited selection if that were the problem, he could pick all 5 from brand A

OpenStudy (anonymous):

Yes. I misinterpreted the question, had him selecting pills at random from a mixture of them. Makes little sense in retrospect.